OpenStudy (anonymous):

for the helix r(t) = sin(t)i + cos(t)i + tk (a) find the angle between the helix and the plane at the point of intersection

OpenStudy (anonymous):

\[cos (theta) = \frac{(-i +k)(k)}{|-i +k||k|}\]

OpenStudy (anonymous):

i got up to here

OpenStudy (anonymous):

but im having probvs solving this.

OpenStudy (anonymous):

i just just need to solve that eqn above dw about the question

OpenStudy (anonymous):

i forgot how to solve cos(theta) blah blah

OpenStudy (anonymous):

the answer is \[\frac{1}{sqrt(2)}\]

OpenStudy (anonymous):

but how do i get that???

ganeshie8 (ganeshie8):

where is the plane ?

OpenStudy (anonymous):

normal to the plane is k

OpenStudy (anonymous):

i have already done that step

OpenStudy (anonymous):

i neeed help with [\cos(theta) = \frac{(-i+k)(k)}{| -i + k||k|}\]

OpenStudy (anonymous):

darn it latex fail

ganeshie8 (ganeshie8):

ahh thats easy then

ganeshie8 (ganeshie8):

|-i+k| = sqrt(1^2 + 1^2) = sqrt(2) |k| = 1

ganeshie8 (ganeshie8):

(-i+k) . (k) = -i.k + k . k = 0 + 1 = 1

OpenStudy (anonymous):

OHHH

ganeshie8 (ganeshie8):

its a dot product ^

OpenStudy (anonymous):

ya OHHHHH smh at the person who supposedly knows everything

OpenStudy (anonymous):

(-i)(k) = 0?

ganeshie8 (ganeshie8):

-i . k = 0 cuz i and k are defined to be perpendicular

ganeshie8 (ganeshie8):